Check whether the value given in the brackets is a solution to the given equation or not:
 
 4p – 3 = 13 (p = – 4)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

LHS = 4p - 3
By substituting the value of p = - 4
Then,
LHS = 4p - 3
= (4 × (-4)) - 3
= -16 - 3
= -19
By comparing LHS and RHS
-19≠13
LHS≠RHS
Hence, the value of p = -4 is not a solution to the given equation 4p - 3 = 13

Thanks